OpenStudy (anonymous):

The surface area A of a bird’s wings is dependent on its weight W in the formula A = 100 to the third power times w to the second power a) Find the area of the wings if the weight is 8 pounds. b) Write the formula with rational exponents.

OpenStudy (anonymous):

\[a = 100^{3}\sqrt{8^{2}}\]

OpenStudy (anonymous):

\[\sqrt{2r ^{5}}*\sqrt{10t ^{4}}\]

OpenStudy (nottim):

i wouldnt kno srry